OpenStudy (anonymous):

A cube is roled 360 times and the results are recorded as: 96 ones, 31 two's, 49 three's, 76 fours, 45 fives, and 63 sixes. What is the experimental probability of rolling a 2 or a 3? PLEASE ANSWER!

OpenStudy (anonymous):

80.

OpenStudy (kropot72):

The experimental probability of rolling a 2 or a 3 is given by: \[\frac{31+49}{360}=\frac{80}{360}=\frac{2}{9}\]

OpenStudy (anonymous):

Thank you

OpenStudy (kropot72):

You're welcome :)